OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work: -7/12 + (-3/5)

OpenStudy (anonymous):

= -7/12 + (-3/5) ....... Your expression = -7/12 -3/5 ....... + * - = - (when you multiply a positive number with negative the answer is negative) = -(7/12 + 3/5) ....... taking negative common = -(7*5 + 12*3/12*5) ...... Adding fractions = -(35+36/60) ....... Simplifying = -(71/60) = -71/60 ..... Your Answer

OpenStudy (madhu.mukherjee.946):

@gkidus nice job:)

OpenStudy (anonymous):

I have another question will the answer change if the 7/12 was not a negative.

OpenStudy (michele_laino):

of course!

OpenStudy (anonymous):

So it would just make the answer a positive or will it be a whole different answer

OpenStudy (michele_laino):

in that case we have this: \[\Large \frac{7}{{12}} - \frac{3}{5} = \frac{{\left( {7 \cdot 5} \right) - \left( {3 \cdot 12} \right)}}{{60}} = ...?\]

OpenStudy (anonymous):

So 7/12+(-3/5) would be worked out like this:7/12-3/5=(7•5)-(3•12)/60 =

OpenStudy (yolomcswagginsggg):

7 12 − 3 5 ? = (7)(5)− ((3)(12)) 60 −0.016667≠34.4